OpenStudy (anonymous):

the graphs of the equations y=2x-7 and y-ky=7 are parallel when k equals..?

OpenStudy (anonymous):

two lines are parallel if their slopes are the same

OpenStudy (anonymous):

so how do i solve this?

OpenStudy (anonymous):

equation of a line in the slope intercept form is y = mx+b. where m is the slope.

OpenStudy (anonymous):

so not the first equation but the second one, i have to put it in y=mx+b right?

OpenStudy (anonymous):

yes, put the second equation in the slope intercept form.

OpenStudy (anonymous):

how would i do that

OpenStudy (anonymous):

wait. are you sure the second equation is y - ky = 7?

OpenStudy (anonymous):

noo its y-kx=7

OpenStudy (anonymous):

okay. so express it in the form y = mx+b, in otherwords, isolate y on the LHS and take the x term and constant term over to the RHS

OpenStudy (anonymous):

Y=KX+?

OpenStudy (anonymous):

Y=KX+7?*

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

so the equations y = 2x-7 and y = kx+7 are parallel when their slopes are parallel. for an equation y = mx+b, m is the slope. these hints should be more than enough to help you solve this problem

OpenStudy (anonymous):

sorry I meant to say they are parallel when their slopes are equal.

OpenStudy (anonymous):

its okay, thankyou

OpenStudy (anonymous):

you are welcome.